OpenStudy (anonymous):

Solve for (0,2pi): 2cos^2x+3sinx-3=0

OpenStudy (anonymous):

Trigonometric equations suck .. please help .. :(

OpenStudy (shamim):

plz c my attachment

OpenStudy (anonymous):

Thank you! such great help! :) much appreciated! @shamim

OpenStudy (shamim):

u r most welcome

OpenStudy (shamim):

one more thing is sin x=sin(π/6)=sin(π-π/6)=sin(5π/6) so x=5π/6

OpenStudy (shamim):

if u r pleased on me then u can give me medal

OpenStudy (anonymous):

just one question @shamim when you factored out the negative where did it go? it just disappeare on the 4th step?

OpenStudy (shamim):

-3+2=-1

OpenStudy (shamim):

i used it in my 4th step

OpenStudy (shamim):

still not clear? tell me plz

OpenStudy (anonymous):

yea first you factored out the negative in -2sin^2x+3sinx-1=0 and made it -(2sin^2x-3sinx+1)=0 and then when you rewrote it you put 2sin^2x-3sinx+1=0. Where did the factored out negative go?

OpenStudy (anonymous):

@shamim

OpenStudy (shamim):

i multiplied by a negative both side

OpenStudy (shamim):

u know -0=0

OpenStudy (anonymous):

ohh okay that i get it :) thanks!

OpenStudy (shamim):

u r most welcome